OpenStudy (anonymous):

Find x on a triangle that has a hypotenuse of x, a side of 15 and a side of 20

Directrix (directrix):

You can crank out the value of the hypotenuse using the Pythagorean Theorem. If you want to do that, x^2 = 15^2 + 20^2 and solve for x.

Directrix (directrix):

This problem is "ripe" for application of the Pythagorean Triples. (3,4,5) is one of the infinitely such triples. I know 3^2 + 4^2 = 5^2. In the given triangle, the sides are 15, 20, and x. 15 = 3*5 20 = 4*5 Therefore, x = 5*5 and the hypotenuse is 25. More on Pythagorean Triples at the link: http://www.mathsisfun.com/pythagorean_triples.html